Write the absolute value equations in the form |x - b| = c that have the following solution sets.
a) One Solution: x = -3.
b) One Solution: x = -4.1

Answer by ikleyn(52835)

.

(a)  The equation  is   |x - (-3)| = 0,


     or equivalently    | x + 3  | = 0.


     b = -3;  c = 0.





(b)  The equation   is  |x - (-4.1)| = 0,


     or equivalently    | x + 4.1  | = 0.


     b = -4.1;  c = 0.

--------------

Solved, answered and explained.